OpenStudy (anonymous):

how do I plot \[\left| z - i \right| \ge 2\] ? I already formed it to sqrt{x² + (y-1)²} \[\ge2\] or rather sqrt{x² + y² - 2y +1} \[\ge2\] but what now?

OpenStudy (praxer):

then square both sides, the left hand side Sqrt. shall be removed and see if it works... :)

OpenStudy (praxer):

I think you will get $$x^2+y^2-2y \ge 3$$ we can write this as, $$(x-0)^2+(y-1)^2=4 \ or 2^2 $$ this is the equation of a circle. I think, this is what you require. :)

OpenStudy (praxer):

https://www.desmos.com/calculator/psgdb2yvmx

OpenStudy (praxer):

Did you understand this @snuggles7. :)

OpenStudy (anonymous):

are you a wizard? o.o how do I get (x−0)²+(y−1)²=4 or2² from x²+y²−2y≥3

OpenStudy (anonymous):

I've heard of the circle equation but I never actually used it @__@"

OpenStudy (anonymous):

aa...! I fathom this is a case of 'forcing a formula' on a mathematical problem like expanding for the sake of obtaining the 3rd binomial formula! thank you, yes I understood

OpenStudy (anonymous):

buuuu~t... why does wolfram alpha show me THIS: http://www.wolframalpha.com/input/?i=plot+ |z-i|+%3E+2

OpenStudy (praxer):

https://www.desmos.com/calculator/psgdb2yvmx